# Flexible-length Text Infilling for Discrete Diffusion Models

### v1

@@ Appendix A Appendix
This appendix provides supplementary material to support the findings and methodology presented in the main paper. It includes detailed model size comparisons, generation diversity analysis, hyperparameter sensitivity studies, visualizations of token path dynamics, and the DDOT training algorithm.
A.1 Size Comparison
Table 6 presents a detailed breakdown of performance metrics (BLEU-2/4, NIST-2/4, and METEOR) for different sizes of our DDOT model (small and medium) and its variants (DDOT-R and DDOT-U). These results are shown for the One-Billion-Word-Random and Yelp-Random benchmarks. This comparison allows for an assessment of model scalability and the impact of the $z_{T}$ initialization strategy (random vs. uniform) across different model capacities. The medium-sized DDOT-U variant generally demonstrates the strongest performance, highlighting the benefits of uniform initialization and increased model parameters.
@@ A.1 Size Comparison
A.2 Generation Diversity
Table 7 provides an evaluation of generation diversity for DDOT and baseline models on the One-Billion-Word-Random and Yelp-Random datasets. Diversity is measured using D2 and D4 metrics, which represent the percentage of unique bi-grams and 4-grams in the generated outputs, respectively. These metrics are crucial for assessing the variety and richness of the generated text beyond simple n-gram overlap with references. The results show how DDOT variants compare to other baselines like CBART, LC, and PoP in this regard, with CBART exhibiting high diversity scores. DDOT models offer competitive diversity, particularly DDOT-U on the One-Billion-Word dataset.
@@ A.2 Generation Diversity
A.3 Hyperparameter Sensitivity Analysis
This section details the hyperparameters explored and used in our experiments, providing transparency and aiding reproducibility.
A.3.1 $\lambda$ values
The choice of the loss weight $\lambda$ is critical as it balances the contribution of the token prediction loss ($\mathcal{L}_{\text{tok}}$) and the position prediction loss ($\mathcal{L}_{\text{pos}}$). Table 8 investigates the impact of different $\lambda$ values on model performance metrics (B2, B4, N2, N4, M) for the One-Billion-Word dataset. The results demonstrate the model’s performance across a range of $\lambda$ values, indicating a degree of robustness. For the experiments reported in this table, a $\lambda$ value around 10 appears to offer a good balance.
@@ A.3.2 Training Hyperparameters
Table 9 lists the primary hyperparameters used for training the DDOT models presented in the main paper. These settings, including batch size variations (dependent on GPU resources like A30, L40, or A100), the $\lambda$ value used with scaling (300), weight decay, learning rate, optimizer choice (AdamW), and its $\beta$ parameters, are provided for reproducibility. These parameters were selected based on preliminary experiments and common practices in training diffusion models. Most parameters were selected based on SEDD’s values. $\lambda$ was emperically determined from small test runs.
@@ A.4 Path Visualizations
The following figures provide qualitative insights into the token path dynamics during the diffusion process. These visualizations help to understand the behavior of token positions during inference (sampling) and the nature of the ground truth paths the model learns during training, both with and without Optimal Transport (OT) coupling.
Figure 6 visualizes token trajectories during the inference (sampling) process without OT coupling, shown for both scaled (a) and unscaled (b) positions across a number of sampling steps. These plots illustrate how token paths can become less orderly and may curve or cross significantly towards the end of the generation process when OT is not employed to maintain relative order, potentially leading to permutation errors.
(a) Scaled positions over time without OT
(b) Unscaled positions over time without OT
@@
Figure 5: Ablations on number of sampling steps. Token paths tend to curve towards the end of generation without OT coupling.
Figure 6, in contrast, shows token trajectories during inference (sampling) *with* OT coupling, again for scaled (a) and unscaled (b) positions. These paths are visibly straighter and more ordered throughout the generation process. This demonstrates the effectiveness of OT in maintaining relative token order and guiding tokens towards their target positions more directly, leading to more stable and accurate generation.
(a) Scaled positions over time with OT
(b) Unscaled positions over time with OT
Figure 6: Ablations on number of sampling steps. Token paths tend to be straighter towards the end of generation with OT coupling.
Figure 7 depicts the ground truth token paths between $z_{T}$ (initial noisy state) and $z_{0}$ (target state) during training, scaled between -1 and 1. Subfigure (a) shows these paths *without* OT coupling; here, $z_{T}$ represents a more arbitrary mapping from initial random positions to $z_{0}$, leading to many crossing paths. This illustrates the complex, potentially tangled mapping the model would need to learn. Subfigure (b) shows the paths *with* our sample-level OT coupling. In this case, $z_{T}$ is aligned with $z_{0}$ via optimal transport (within prompt and response sets), resulting in significantly straighter, non-crossing (within-set) trajectories. This comparison highlights how OT simplifies the learning objective for the positional diffusion component by providing more direct and disentangled target paths.
(a) Scaled positions over time without OT
(b) Scaled positions over time with OT
Figure 7: Ground Truth Token Velocities (scaled between -1 and 1). The ground truth velocities without OT (a) have a lot of crossing lines demonstrating instability in matching whereas the velocities with OT coupling (b) show almost straighter lines throughout the denoising process.
A.5 DDOT Training Algorithm
Algorithm 1 provides a concise overview of the DDOT training procedure. It details the steps involved in each training iteration:
  1. 1.
    Sampling a data instance $(x_{0},z_{0})$ from the dataset $\mathcal{D}$.
  2. 2.
    Sampling initial noisy positions $z_{T}\sim\mathcal{U}(-1,1)^{L}$ and a random timestep $t\sim\mathcal{U}(0,1)$.
  3. 3.
    Permuting the sampled $z_{T}$ to minimize the Optimal Transport cost with respect to the target positions $z_{0}$. This step is crucial for establishing the non-crossing paths within prompt and response segments.
  4. 4.
    Linearly interpolating between $z_{0}$ and the permuted $z_{T}$ to obtain the positions $z_{t}$ for the sampled timestep $t$.
  5. 5.
    Calculating the total loss $\mathcal{L}$ as a weighted sum of the token prediction loss $\mathcal{L}_{\text{tok}}$ and the position prediction loss $\mathcal{L}_{\text{pos}}$.
  6. 6.
    Updating the model parameters $\theta$ (which includes parameters for both $s_{\theta}$ and $v_{\theta}$) using the gradient of the total loss $\nabla\mathcal{L}$.
This process is repeated until convergence.
Algorithm 1 DDOT Training
1:Dataset $\mathcal{D}$, model $s_{\theta},v_{\theta}$, loss weight $\lambda$
2:while Training do
3:     $(x_{0},z_{0})\sim\mathcal{D}$; $z_{T}\sim\mathcal{U}(-1,1)^{L}$; $t\sim\mathcal{U}(0,1)$
4:     Permute $z_{T}$ to minimize OT cost
5:     Interpolate positions $z_{t}=(1-t)z_{0}+tz_{T}$
6:     $\mathcal{L}\leftarrow\mathcal{L}_{\text{tok}}+\lambda\mathcal{L}_{\text{pos}}$
7:     $\theta\leftarrow\text{Update}(\theta,\nabla\mathcal{L})$
8:end while

### v2

@@ Appendix A Appendix
A.1 Overview
This appendix provides additional results, diversity analyses, hyperparameter sensitivity, path visualizations, and short implementation notes for reproducibility. We reuse notation from the main text and make one explicit addition used below:
$\ell^{\mathrm{np}}\;=\;L-\ell^{p}$
denoting the number of non-prompt slots (response ++ pads). This matches the usage of $z_{T}^{\mathrm{np}}$ in the main text.
A.2 Additional Results
A.2.1 Size Comparison
Table 6 reports BLEU-2/4, NIST-2/4, and METEOR for ddot small/medium and for ddot-R vs. ddot-U on One-Billion-Word–Random and Yelp–Random. Uniform $z_{T}$ initialization (ddot-U) consistently outperforms random (ddot-R), and scaling to the medium model further improves most metrics.
@@ A.2.1 Size Comparison
A.2.2 Generation Diversity
Table 7 compares diversity using D2/D4 (unique bigrams/four-grams). ddot-U yields competitive diversity relative to strong NAR baselines while maintaining ordering constraints.
@@ A.2.2 Generation Diversity
A.3 Hyperparameter Sensitivity and Training Details
A.3.1 Loss Weight $\lambda$
We sweep $\lambda$ to balance $\mathcal{L}_{\text{tok}}$ and $\mathcal{L}_{\text{pos}}$. Results in Table 8 show stable performance across a range; $\lambda\!\approx\!10$ is a good default used in the main results.
@@ A.3.2 Training Hyperparameters
[…]
@@ A.4 Path Visualizations
We visualize token-position trajectories during both inference (what the model actually does at test time) and training (the straight paths used as supervision). We show two coordinate systems: scaled to [1,1][-1,1] (used in losses; cf. Equation 3) and unscaled in the model’s native context length (for intuition about absolute movement). Unless noted, we split by set $S\!\in\!\{p,r\}$ and index tokens within a set by their ground-truth order.
Plot Explanation.
Let $\{t_{u}\}_{u=0}^{K}$ be the reverse-time grid (left-to-right on the $x$-axis), with $t_{0}=T$ and $t_{K}=0$. For each token index $i$ in set $S$, we plot its 1D position $z^{S,(i)}_{t_{u}}$ across $u$. Colors denote the token’s state (masked vs. unmasked) at each step. During inference, $z_{t_{u+1}}$ is obtained by an Euler update using the learned velocity $v_{\theta}(z_{t_{u}},x_{t_{u}},t_{u})$ while clamping prompts and keeping pads stationary (Equation 5). During training, targets follow straight paths $z_{t}^{S}=(1-t)z_{0}^{S}+t\,\tilde{z}_{T}^{S}$ from Equation 8.
Analysis.
Within a set (prompt or response), desirable behavior is (i) order preservation (no within-set crossings) and (ii) near-straight trajectories (little curvature). Inter-set crossings (prompt vs. response) are allowed—and often required—to place response tokens between prompts during infilling.
(a) Scaled positions over time without OT.
(b) Unscaled positions over time without OT.
@@
Figure 5: Ablations on number of sampling steps without OT coupling. Paths curve and cross more often, complicating learning and inference.
(a) Scaled positions over time with OT.
(b) Unscaled positions over time with OT.
Figure 6: With per-set OT coupling, trajectories are straighter and order-preserving within each set.
(a) Scaled training targets without OT: many crossings.
(b) Scaled training targets with OT: straighter, non-crossing (within-set).
Figure 7: Ground-truth token paths (scaled to [1,1][-1,1]) used as supervision. Without OT (left), many crossings indicate unstable matching; with OT (right), trajectories are nearly straight and order-preserving within each set.
A.5 Implementation Notes
Computing OT efficiently. Prompt matching is solved exactly in 1D by sorting $z_{0}^{p}$ and $z_{T}^{p}$ and pairing by rank; response matching uses order-preserving injective alignment via global-alignment DP against $z_{T}^{\mathrm{np}}$. Both avoid dense cost matrices and run in $O(\ell^{p}\log\ell^{p}+\ell^{r}(L-\ell^{p}))$ time and $O(L)$ memory.
Caching & streaming. For large corpora, precompute couplings on CPU and stream mini-batches; we observed negligible overhead whether caching or computing couplings on the fly.